OpenStudy (superhelp101):

Let f(x) = 2x+3 and g(x) = x2 Part 1 [2 points] Find f(g(x)) Part 2 [2 points] Find g(f(x)) Part 3 [4 points] Use complete sentences to explain any difference in part 1 and part 2.

OpenStudy (superhelp101):

@Hero @mathmale @mathstudent55

hero (hero):

f(x) = 2x + 3 g(x) = x^2 f(g(x)) = 2(g(x)) + 3 g(f(x)) = f(x)^2

OpenStudy (superhelp101):

ok thanks! can you help me on part 3? :)

hero (hero):

For f(g(x)), the entire function g(x) is used as the input into f(x). For g(f(x)), the entire function f(x) is used as the input into g(x).

hero (hero):

For Part 1 and 2, You still have to figure out what 2(g(x)) + 3 is equal to f(x)^2 is equal to The final expression should be completely algebraic.

OpenStudy (superhelp101):

how do I figure that out?

hero (hero):

Insert the original expression for g(x) into 2(g(x)) + 3 then simplify. Insert the original expression for f(x) into f(x)^2 then simplify.

OpenStudy (superhelp101):

oh so it would like this: ....

OpenStudy (superhelp101):